OpenStudy (anonymous):

(x)(1/15)+(x+1)(1/20)=1

OpenStudy (bahrom7893):

wait im doing this wrong..

OpenStudy (bahrom7893):

x/15 + x/20 + 1/20 = 1 20x/15 + x + 1 = 20 20x/15 +x = 19

OpenStudy (bahrom7893):

(20x + 15x)/15 = 19

OpenStudy (bahrom7893):

35x/15 = 19 7x/3 = 19 7x = 19*3 x = 8.142

OpenStudy (bahrom7893):

satellite check my arithmetic haha..

OpenStudy (anonymous):

you can also multiply both sides by 60 to clear the fractions you will get \[4x+3(x+1)=60\][ \[7x+3=60\] \[7x=57\] \[x=\frac{57}{7}\]

OpenStudy (anonymous):

is that the same answer?

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

ok then it is right. but i like getting rid of fractions to make the arithmetic easier. of course it doesn't matter.